OpenStudy (anonymous):

find the range of k if the quadratic equation 2x^2-x=k has real and distinct roots.

OpenStudy (amistre64):

-b/2a will give you the lowest point

OpenStudy (amistre64):

or rather, when x= -b/2a thats the one lol

OpenStudy (anonymous):

so the answer is?

OpenStudy (amistre64):

that is the answer .....

OpenStudy (anonymous):

oh god, just do b^2 - 4ac. (-1)^2 - (4 * 2 * -k) < 0 so 1 + 8k < 0 so k< (-1/8)

OpenStudy (anonymous):

lol okay thanks :)

OpenStudy (anonymous):

@tanvidais, shoudn't it be b^2-4ac = 0 ?

OpenStudy (anonymous):

because it has real and distinct roots right?

OpenStudy (anonymous):

um if the discriminant is <0 it has no real roots ,=0 has 1 real root, >0 has 2 real roots i thought

OpenStudy (anonymous):

and it says it has real and distinct roots. so the discriminant supposed to be =0 right?

OpenStudy (amistre64):

the range of: 2x^2-x = k is no different from the range of: 2x^2-x=y 2x^2-x = k 2(x^2-1/2 x) = k 2(x^2-1/2 x + 1/16) - 1/8 = k 2(x-1/4)^2 - 1/8 = k it looks to be then from -1/8 to infinity if im reading it right

OpenStudy (anonymous):

oh, shoot. then it's k> (-1/8)

OpenStudy (anonymous):

>0 since it has 2 diff roots it says in question "real and distinct"

OpenStudy (anonymous):

when it is real roots, it is => 0. when it is real and distinct, it is > 0. when it is equal roots, it is = 0. when it is no roots, it is <0.

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

no problem :)

OpenStudy (amistre64):

2x^2 -x -k ; such that k > -1/8 has 2 real and disticnt roots